# Electronic Structure of Epitaxial Films of the Bilayer Strontium Ruthenate: Sr3Ru2O7

### v1

@@ Abstract
We report the first combined study of the low-energy electronic band structure of epitaxial Sr3Ru2O7 films using angle-resolved photoemission spectroscopy (ARPES) and density functional theory (DFT). The complete Fermi-surface topography of the near-Fermi-level bands is determined from in-situ ARPES measurements. To investigate the effects of substrate-induced strain on the band structure, Sr3Ru2O7 thin films are epitaxially grown on SrTiO3 (STO) and (LaAlO3)0.3(Sr2TaAlO6)0.7 (LSAT) substrates using molecular beam epitaxy. The combination of the measured Fermi-surfaces along with the theoretical interpretation, clearly show dramatic changes in the Fermi surface topologies that result from the underlying strain states of the films on the two substrates. We find that the Sr3Ru2O7 films prepared on STO are tensile strained with tetragonal symmetry, whereas those grown on LSAT are compressively strained with orthorhombic symmetry. Within below the Fermi level, we observe two flat bands along - in the orthorhombic phase and around in the tetragonal phase. These features could be favorable for van Hove singularities near the Fermi level, and highlight the emergence of magnetic instabilities in epitaxial Sr3Ru2O7 films.
@@
Any modification of the Fermi surface (FS) induced by strain, whether subtle or pronounced, is expected to influence the dynamics of magnetic-field-controlled instabilities and may give rise to new phases [27, 1, 3, 38]. Changes in the FS topology arise from alterations of the underlying crystal symmetry, allowing strain engineering to provide precise control over the electronic and magnetic interactions in bilayered Sr3Ru2O7 [62, 55, 11]. This tunability offers a powerful platform for exploring electronic and magnetic ordered phases, quantum criticality, and their relationship to the underlying band structure [58, 11, 38].
@@
Here, we report the first combined experimental and theoretical investigation of the electronic band structure and FS topology of Sr3Ru2O7 thin films. Using molecular beam epitaxy (MBE), we deposited Sr3Ru2O7 films on LSAT and STO single-crystal substrates. The reflection high-energy electron diffraction (RHEED) data acquired during growth indicate high-quality epitaxial Sr3Ru2O7 films; while the in-plane lattice parameters extracted from reciprocal space maps (RSMs) confirm coherent growth accompanied by compressive and tensile strains on LSAT and STO, respectively. To examine how these strain states modify the electronic structure, we performed ARPES measurements on the Sr3Ru2O7 films. We observe substantial strain-dependent transformations in the band dispersions. The measured FSs show that compressively strained films adopt an orthorhombic structure, whereas tensile-strained films exhibit a tetragonal symmetry. These observations indicate that the RuO6 octahedral rotation responsible for Brillouin zone reconstruction and band backfolding is suppressed in samples grown on STO. The DFT-calculated bands, obtained by incorporating the RSM-derived lattice parameters, are in general agreement with the ARPES experiments and reproduce the main features of the measured FSs. Within a few meV below EF, we observe two flat bands along - in the orthorhombic phase and around in the tetragonal phase. These bands suggest possible van Hove singularities near EF, a condition that favors magnetic instabilities in epitaxial Sr3Ru2O7 films.
@@
In-situ ARPES experiments were conducted at the temperature of 7 K, using a state-of-the art high-resolution Scienta DA-30L electron analyzer with an energy resolution of 5 meV. Figure 1(b) shows the angle-resolved geometry of the ARPES experiment. For these measurements, 6 nm thick Sr3Ru2O7 samples were used. This was the optimal thickness ensuring high surface quality for the acquisition of high quality ARPES data. The theoretical investigations were carried out using pseudopotential, plane-wave implementation of DFT as described in Ref. [40, 13].
The growth of the Sr3Ru2O7 thin films was monitored in real time using RHEED. The diffraction patterns of the films grown on LSAT and STO are shown in Fig. 1(c) and 1(d), respectively. These patterns indicate good film quality and the epitaxial growth. In particular, the presence of the very sharp diffraction spots attests to the two dimensional growth mode of the Sr3Ru2O7 films on both substrates, as well as to the flatness, the smoothness, and the perfect single crystalline structure of their surfaces [22, 43, 7].
@@
The ARPES-measured FSs reveal substrate-dependent modifications that reflect the underlying strain states of the films. The tensile strain imposed by the STO substrate largely preserves tetragonal symmetry, whereas the compressive strain from the LSAT substrate reduces the symmetry and induces an orthorhombic phase. These structural differences, including the resulting fourfold rotation symmetries around , are well captured by the DFT calculations. Notably, while orthorhombic systems do not strictly possess fourfold symmetry, our observations remain consistent with previous ARPES and DFT studies on Sr3Ru2O7 single crystals [55, 47], indicating that the bilayer and substrate constraints dominate the observed electronic structure. Another important consideration is the low temperature at which the ARPES measurements are performed, as both LSAT and STO substrates are known to undergo symmetry changes that could, in principle, influence the film structure in addition to strain effects [6, 25]. However, the presence of orthorhombic and tetragonal symmetries at room temperature can be directly inferred from reciprocal space mapping (RSM) measurements, reducing the likelihood that the observed structural changes arise from temperature-driven substrate transitions or interfacial coupling effects. To further corroborate the strain-driven structural distortions inferred from ARPES and DFT, the octahedral rotation patterns can be identified through half-integer reflections measured using synchrotron-based XRD, which are particularly sensitive to strain states and interfacial coupling [15, 21, 59, 31, 32]. Such measurements, however, are beyond the scope of the present study.

### v2

@@ Abstract
We report the first combined study of the low-energy electronic band structure of epitaxial Sr3Ru2O7 films using angle-resolved photoemission spectroscopy and density functional theory. The complete Fermi-surface topography of the near-Fermi-level bands is determined from in-situ ARPES measurements. To investigate the effects of substrate-induced strain on the band structure, Sr3Ru2O7 thin films are epitaxially grown on SrTiO3 and (LaAlO3)0.3(Sr2TaAlO6)0.7 substrates using molecular beam epitaxy. The combination of the measured Fermi-surfaces along with the theoretical interpretation, clearly show dramatic changes in the Fermi surface topologies that result from the underlying strain states of the films on the two substrates. We find that the Sr3Ru2O7 films prepared on SrTiO3 are tensile strained with tetragonal symmetry, whereas those grown on (LaAlO3)0.3(Sr2TaAlO6)0.7 are compressively strained with orthorhombic symmetry. Within below the Fermi level, we observe two flat bands along - in the orthorhombic phase and around in the tetragonal phase. These features could be favorable for van Hove singularities near the Fermi level, and highlight the emergence of magnetic instabilities in epitaxial Sr3Ru2O7 films.
@@
Any modification of the FS induced by strain, whether subtle or pronounced, is expected to influence the dynamics of magnetic-field-controlled instabilities and may give rise to new phases [27, 1, 3, 38]. Changes in the FS topology arise from alterations of the underlying crystal symmetry, allowing strain engineering to provide precise control over the electronic and magnetic interactions in bilayered Sr3Ru2O7 [62, 55, 11]. This tunability offers a powerful platform for exploring electronic and magnetic ordered phases, quantum criticality, and their relationship to the underlying band structure [58, 11, 38].
@@
Here, we report the first combined experimental and theoretical investigation of the electronic band structure and FS topology of Sr3Ru2O7 thin films. Using molecular beam epitaxy (MBE), we deposited Sr3Ru2O7 films on (LaAlO3)0.3(Sr2TaAlO6)0.7 (LSAT) and SrTiO3 (STO) single-crystal substrates. The reflection high-energy electron diffraction (RHEED) data acquired during growth indicate high-quality epitaxial Sr3Ru2O7 films; while the in-plane lattice parameters extracted from reciprocal space maps (RSMs) confirm coherent growth accompanied by compressive and tensile strains on LSAT and STO, respectively. To examine how these strain states modify the electronic structure, we performed ARPES measurements on the Sr3Ru2O7 films. We observe substantial strain-dependent transformations in the band dispersions. The measured FSs show that compressively strained films adopt an orthorhombic structure, whereas tensile-strained films exhibit a tetragonal symmetry. These observations indicate that the RuO6 octahedral rotation responsible for Brillouin zone reconstruction and band backfolding is suppressed in samples grown on STO. The DFT-calculated bands, obtained by incorporating the RSM-derived lattice parameters, are in general agreement with the ARPES experiments and reproduce the main features of the measured FSs. Within a few meV below EF, we observe two flat bands along - in the orthorhombic phase and around in the tetragonal phase. These bands suggest possible van Hove singularities near EF, a condition that favors magnetic instabilities in epitaxial Sr3Ru2O7 films.
@@
In-situ ARPES experiments were conducted at the temperature of 7 K, using a state-of-the art high-resolution Scienta DA-30L electron analyzer with an energy resolution of 5 meV. Figure 1(b) schematically illustrates the angle-resolved geometry of the ARPES experiment. For these measurements, 6 nm thick Sr3Ru2O7 samples were used. This was the optimal thickness ensuring high surface quality for the acquisition of high quality ARPES data. The theoretical investigations were carried out using pseudopotential, plane-wave implementation of DFT as described in Ref. [40, 13].
The growth of the Sr3Ru2O7 thin films was monitored in real time using RHEED. The diffraction patterns of the films grown on STO and LSAT are shown in Fig. 1(c) and 1(d), respectively. These patterns indicate good film quality and the epitaxial growth. In particular, the presence of the very sharp diffraction spots attests to the two dimensional growth mode of the Sr3Ru2O7 films on both substrates, as well as to the flatness, the smoothness, and the perfect single crystalline structure of their surfaces [22, 43, 7].
@@
The ARPES-measured FSs reveal substrate-dependent modifications that reflect the underlying strain states of the films. The tensile strain imposed by the STO substrate largely preserves tetragonal symmetry, whereas the compressive strain from the LSAT substrate reduces the symmetry and induces an orthorhombic phase. These structural differences, including the resulting fourfold rotation symmetries around , are well captured by the DFT calculations. Notably, while orthorhombic systems do not strictly possess fourfold symmetry, our observations remain consistent with previous ARPES and DFT studies on Sr3Ru2O7 single crystals [55, 47], indicating that the bilayer and substrate constraints dominate the observed electronic structure. Another important consideration is the low temperature at which the ARPES measurements are performed, as both LSAT and STO substrates are known to undergo symmetry changes that could, in principle, influence the film structure in addition to strain effects [6, 25]. However, the presence of orthorhombic and tetragonal symmetries at room temperature can be directly inferred from RSM measurements, reducing the likelihood that the observed structural changes arise from temperature-driven substrate transitions or interfacial coupling effects. To further corroborate the strain-driven structural distortions inferred from ARPES and DFT, the octahedral rotation patterns can be identified through half-integer reflections measured using synchrotron-based XRD, which are particularly sensitive to strain states and interfacial coupling [15, 21, 59, 31, 32]. Such measurements, however, are beyond the scope of the present study.